Use the method of variation of parameters to find the general solution of the system

x = [2 21]x+[287] Ax + g(t)

Find the Laplace transform

f(t) = S(t – 1)cos (t)
Answer #1

Given x = [e &] x + ( 2ct - Ax+ g(t) Variation of parameters - for general solution XG=XH*Xp - for homogeneous solution, get)for 2 2 = 1 2 -- -] A-C M-a ( 2x3 = [] 2,=82 - 2x, +2 Xe=0 - 2 2x - 2Xq=0 X = X2 take x2=1 21=-3 - 2 = 1 vp = { homogeneous sxp (t) = X(t) X(t) & (t) et cest et - est e3t lt et ét 3t - e 3t ot -2e²t + 3te3t dt 2 e3t et 2e2t +3tét -est ezt + 3t 9 ca 3we know that 6 Given L ( sct-1) cest) [ f(t) SCt-a) = 0 itt ta Jef (t) s(t-a) = f(a) it t= a » Caplace toans form of f (+) {C